# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 5.2.1 Persistent homology of simplicial complexes
We assume in the following that the reader is familiar with simplicial complexes, boundary operators and homology groups, and we refer the interested reader to [Mun84, Chapter 1] for a thorough treatment of these notions. The first important definition is the one of filtered simplicial chain complexes.
Definition 5.3.
Let $S$ be a simplicial complex, and $f\colon S\to\mathbb{R}$ be a filtration function, i.e., $f$ satisfies $f(\sigma)\leq f(\tau)$ when $\sigma\subseteq\tau$. Then, the filtered simplicial chain complex $(S,f)$ is defined as $(S,f)=((C_{t})_{t\in\mathbb{R}},\iota)$, where
@@
Note that $f$ can be used to define an order on the simplices of $S=\left\{\sigma_{i}\right\}_{i=0}^{N}$, by using the ordering induced by the filtration values. In other words, we assume in the following that $f(\sigma_{0})\leq f(\sigma_{1})\leq\dots\leq f(\sigma_{N})$. We also slightly abuse notations and define $C_{i}:=\left<\sigma_{0},\dots,\sigma_{i}\right>$ for any $i\in\llbracket 0,N\rrbracket$ and
$(S,f)=\left(C_{0}\overset{\iota_{0}}{\hookrightarrow}C_{1}\overset{\iota_{1}}{\hookrightarrow}\dots\overset{\iota_{N-1}}{\hookrightarrow}C_{N}=\left<S\right>\right).$ (11)
Then, applying the homology functor $H_{*}$ on this filtered simplicial chain complex yields the following (single-parameter) persistence module
@@
Theorem 5.4 (Persistence pairing, [dMV11, Theorem 2.6]).
@@
Note that while the reduced basis $\{\hat{\sigma}_{1},\dots,\hat{\sigma}_{N}\}$ does not need to be unique, the pairing map ${\rm Low}$ is actually independent of that reduced basis (see [EH10, VII.1, Pairing Lemma]).
@@ 5.2.2 Vineyards algorithm and matching
The vineyards algorithm is a method that allows to find reduced chain bases for filtered simplicial complexes whose simplex orderings only differ by a single transposition of consecutive simplices, that we denote by $(i\,i+1)$.
Proposition 5.5 ([CSEM06]).
Let $S=\{\sigma_{1},\dots,\sigma_{N}\}$ be a (filtered) simplicial complex (with filtration function $f$), and let $B=\{\hat{\sigma}_{1},\dots,\hat{\sigma}_{N}\}$ be a corresponding reduced chain basis. Let $\tilde{f}:S\rightarrow\mathbb{R}$ be a filtration function that swaps the simplices at positions $i$ and $i+1$, i.e., that induces a new filtered simplicial complex $\tilde{S}=\{\sigma_{1},\dots,\sigma_{i+1},\sigma_{i},\dots,\sigma_{N}\}$. Note that the swapped basis ${\rm sw}_{i}^{S}(B):=\{\hat{\sigma}_{1},\dots,\hat{\sigma}_{i+1},\hat{\sigma}_{i},\dots,\hat{\sigma}_{N}\}$ might not be a reduced basis for $(\tilde{S},\tilde{f})$, since the pairing map ${\rm Low}$ might not be well-defined anymore. Fortunately, there exists a change of basis, called vineyard update, that turns ${\rm sw}_{i}^{S}(B)$ into a reduced basis $\tilde{B}$ of $(\tilde{S},\tilde{f})$ in $O(N)$ time, and that comes with a bijective map $\sigma_{\mathrm{vine}}:B\rightarrow\tilde{B}$, called vineyard matching, that can be used to define a matching $\sigma_{\mathrm{vine}}:{\mathcal{B}}(f)\to{\mathcal{B}}(\tilde{f})$.
Note that the vineyards matching can be straightforwardly extended to filtration functions whose induced ordering differs by a whole permutation from the one of the initial filtration function by simply decomposing the permutation into elementary transpositions (using, e.g., Coxeter decompositions).
Application to multi-parameter persistent homology.
When the simplices of $S$ are (partially) ordered from a function $f:S\rightarrow\mathbb{R}^{n}$, i.e., one has $f(\tau)\leq_{n}f(\sigma)$ for any $\tau\subseteq\sigma$, then the function $f$ is called a multi-parameter filtration function of $S$. In that case, applying the homology functor also leads to simplicial homology groups $\{H_{d}(C_{i})\}_{i\in I}$, where $I\subseteq\llbracket 1,N\rrbracket^{n}$ is the (partial) ordering associated to $f$, and these groups are connected by morphisms as long as their indices are comparable in $\mathbb{R}^{n}$. These groups and maps are called the multi-parameter persistent homology associated to the multi-filtered simplicial chain complex $(S,f)$. Similarly to the single-parameter case, when $k$ is a field, one can define the multi-parameter persistence module ${\mathbb{M}}$ associated to $(S,f)$ as the family of vector spaces indexed over $\mathbb{R}^{n}$ defined with the identifications $M_{s}:=H_{d}(C)$ where $C={\bf{\rm Vect}}\{\sigma\in S:f(\sigma)\leq s\in\mathbb{R}^{n}\}$.
We now show that the vineyards algorithm yields an exact matching function, in the sense that it commutes with the exact matching function induced by the underlying module ${\mathbb{M}}$ when it is interval decomposable.
Proposition 5.6.
Let ${\mathbb{M}}$ be an interval decomposable multi-parameter persistence module computed from a finite multi-filtered simplicial chain complex $(S,f)$ over $\mathbb{R}^{n}$. Then, for a small enough $\delta>0$, $\sigma_{\mathrm{vine}}$ is exact.

### v4

@@ 6.2.1 Persistent homology of simplicial complexes
We assume in the following that the reader is familiar with simplicial complexes, boundary operators and homology groups, and we refer the interested reader to (53, Chapter 1) for a thorough treatment of these notions. The first important definition is the one of filtered simplicial chain complexes.
Definition 25.
Let $S$ be a simplicial complex, and $f\colon S\to\mathbb{R}$ be a filtration function, i.e., $f$ satisfies $f(\sigma)\leq f(\tau)$ when $\sigma\subseteq\tau$. Then, the filtered simplicial chain complex $(S,f)$ is defined as $(S,f)=((C_{t})_{t\in\mathbb{R}},\iota)$, where:
  1. 1.
@@
Note that $f$ can be used to define an order on the simplices of $S=\left\{\sigma_{i}\right\}_{i=0}^{N}$, by using the ordering induced by the filtration values. In other words, we assume in the following that $f(\sigma_{0})\leq f(\sigma_{1})\leq\dots\leq f(\sigma_{N})$. We also slightly abuse notations and define $C_{i}:=\left<\sigma_{0},\dots,\sigma_{i}\right$> for any $i\in\llbracket 0,N\rrbracket$, and
[…]
Then, applying the homology functor $H_{*}$ on this filtered simplicial chain complex yields the following (single-parameter) persistence module:
@@
Theorem 15 (Persistence pairing, (30, Theorem 2.6)).
@@
Note that while the reduced basis $\{\hat{\sigma}_{1},\dots,\hat{\sigma}_{N}\}$ does not need to be unique, the pairing map $\rm Low$} is actually independent of that reduced basis, see (35, VII.1, Pairing Lemma).
@@ 6.2.2 Vineyards algorithm and matching
The vineyards algorithm (28) is a method that allows to find reduced chain bases for filtered simplicial complexes whose simplex orderings only differ by a single transposition of consecutive simplices, that we denote by (ii+1)(i\,i+1). This algorithm was later generalized to the setup of zigzag persistence modules in (52). This is the setup that we use in our context as follows. We start with some $n$-parameter persistence module $\mathbb{M}$, such that there exists a finite dimensional, generic, $n$-filtered one-critical simplicial chain complex $(S,F)$, satisfying $\mathbb{M}=H_{*}(F)$. [^5] We drop the dependence on $S$ for simplicity. We also fix an order $(\sigma_{1},\dots,\sigma_{N})$ on the simplices of $S$, and use the following notation (in this section only):
$(x_{1},\dots,x_{N}):=\left(F(\sigma_{1}),\dots,F(\sigma_{N})\right),\quad\textnormal{and }\quad(x_{1}^{l},\dots,x_{N}^{l}):=\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}(\sigma_{1}),\dots,{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}(\sigma_{N})\right),$ (20)
the filtration values (in $\mathbb{R}^{n}$) of the simplices of $F$ and $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}$} respectively, for any given positive line $l$. Note that we always have, for any line $l$:
$\forall 1\leq j\leq N,\quad x_{j}\leq_{n}x_{j}^{l}\quad\textnormal{and}\quad\exists 1\leq k\leq n,\quad(x_{j})_{k}=(x_{j}^{l})_{k}.$ (21)
Finally, we consider the map $\mathrm{ord}\colon{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\mapsto\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right)\in\mathfrak{S}_{N}$ that gives the partial order of the simplices of $F$ according to the filtration $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}$}, completed to a total order by the initial order if necessary, i.e., $\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right)$ is the only permutation $\gamma\in\mathfrak{S}_{N}$ satisfying
[…]
Proposition 16.
Consider $l_{1},l_{2}$ two consecutive diagonal lines such that for any diagonal line $l_{1}\leq l\leq l_{2}$, we have either $\gamma=\gamma_{1}$ or $\gamma=\gamma_{2}$, where $\gamma:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right),\,\gamma_{1}:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}\right)$ and $\gamma_{2}:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\right)$. Then, either $\gamma_{1}=\gamma_{2}$ or there exists an integer $1\leq i\leq n$ such that $\gamma_{1}=\gamma_{2}\circ(i\,i+1)$, and the $i$th vineyard update induces a compatible matching between ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{2}}}\right)$.
Remark 11 (More details about the vineyard matching).
More specifically, the matching function given by the vineyards algorithm corresponds to the following. If $(\tau_{1},\dots,\tau_{N})$ is a reduced basis on the first line, then this algorithm provides an updated basis $(\mathrm{vine}_{i}(\tau_{1}),\dots,\mathrm{vine}_{i}(\tau_{N}))$ that is reduced on the second line. The matching function is then given by the following relation: if $b$ (resp. $d$) is the birthpoint (resp. deathpoint) of a bar in the barcode induced by $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}$}, and generated by some $\tau_{j}$ (resp. $\tau_{k}$), then the endpoint $b$ is matched to the endpoint generated by $\mathrm{vine}_{i}(\tau_{j})$ (resp. $\mathrm{vine}_{i}(\tau_{k})$).
Note that a non-trivial bar $\left[{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j}),{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{k})\right)$ may be matched to an empty bar if ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}((\tau_{j}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{k}))$. In particular, such continuous matching functions lead to indicator summands, which, once split up, lead to the desired candidate decomposition with interval summands.